Summary of the 3mm GMVA Session October, 9-15, 2008 ---------------------------------------------------- This session may be only partly successfull due to loss of PdB and bad weather at Pico Veleta. The other stations appeared to have performed well. An ftp fringe check done on Oct. 9 and 10 showed fringes between the stations which participated in this: EB, ON, PV and PdB. Fringes to PdB show large phase noise, indicative of a problem with the overall phase stability (frequency synthesizer?). A second fringe test done a few hours later and after several system checks at PdB, showed the same problem, so it was decided that PdB stops observing in VLBI mode. Some more details: ------------------ Effelsberg: no major problems. Good weather, however high humidity, in particular in the morning. Tau varied between 0.1 and 0.4, with typical values around 0.2. Tsys = 170 K at zenith. An encoder jump in elevation in the last third of the session was discovered immediately and did not affect the VLBI data. Oct. 9 day 283 clear (BLLAC) Oct. 10 day 284 scattered clouds, afternoon clear (MrK421) Oct. 11 day 285 foggy in the morning, later thin clouds (3C84, 0716+714) Oct. 12 day 286 foggy in the morning, later clear (Marscher project, OJ287) Oct. 13 day 287 thin clouds in the morning, very cloudy onm the afternoon (0528+13, OJ287) Oct. 14 day 288 first cloudy, later blue sky (0716+714, NRAO150) Oct. 15 day 289 thin clouds, still good (3C111) Pico Veleta: high wind, bad weather during c082a, almost no observations done weather still bad, but some observations done for c082b day observed 9oct 10% 21:15 to 22:22 10oct 0% 11oct 0% 12oct 25% 13:34 to 19:52 13oct 50% 08:40 to 17:37; 19:45 to 23:52 14oct 60% 00:00 to 01:52; 05:00 to 10:22; 16:30 to 23:52 15oct 100% 00:00 to 10:07 total lost time about 65% due to wind, snow, rain but even when observing the weather was not good most of the time. PdB: excessive phase noise discovered on Oct. 9 after ftp-fringe check. Station went back to local observing in non-VLBI mode. Onsala: all projects observed. Weather was not really great, sometimes even rainy. Metsahovi: receiver problems, RCP did not work. Later during the experiment, and we don't know at which point, the carousel from the Russian receiver got stuck. That means that the beam of the receiver wasn't really on the sky. In addition, we had a power cut during the Monday morning which provoked loss of scans. As we are not sure if the data on the diskpacks is valuable, do you want to take a look yourself to the log files and decide if to ship them or instead keep them here for the next session? I know that the news about the experiment are not really promising, but we are hoping to get asap the new receiver so we can work properly in the next sessions. VLBA: Weather from daily webcam inspection Oct. 9 FD cloudy BR,OV,KP,LA,PT,NL almost clear Oct. 10 all stations clear Oct. 11 BR,KP,OV clear NL,FD,LA,PT,MK scattered light clouds afternoon: PT, KP, LA rain Oct. 12 LA overcast NL,FD,PT,KP,OV some clouds MK, BR clear Oct. 13 BR,OV,FD,PT,KP,MK clear NL overcast LA broken web-cam ? Oct. 14 KP,OV,MK,PT clear FD,LA rain NL,BR cirrus Oct. 15 NL,FD,LA rain PT,KP,OV clear BR,MK thin clouds VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C082A Observer: Krichbaum, T. Project type: VLBI Obs filename: c082acrd.* Date/Day: 2008OCT09/283 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operators are Heidi Gerhardt and Anthony Sowinski GPS delta, UT1-UTC (-0.50), vclock ok 1856 Begin 1929 *LA 35+ mph wind caution 2059 KP reports wea: clear, dry. 2140 New Operator: Paul Padilla 2302-2304 *KP autostowed, power glitch COM 100 2 0000 Date is 2008OCT10/284. 0014-0023 *OV AZ #1 armature over current ACU 100 9 0046 %FD WEA: Storm front, light rain, wind 0540 New Operator is Alan Kerr 0627-0630 *PT HC2 FIFO error. 21522 100 3 0707-0709 *MK FRM focus drive fault. FRM 100 2 1017 %KP winds are 10m/s with 18m/s gusts. 1031-1034 *KP out of point. COM 100 3 1042-1045 *KP HC2 FIFO. cleared. MCB 100 3 1054-1058 *PT HC2 FIFO error. cleared. 21522 100 4 1123-1128 *PT HC2 FIFO error. cleared. 21522 100 5 1210-1213 *PT HC2 FIFO error. cleared. 21522 100 3 1256-1301 *PT HC2 FIFO error. cleared. 21522 100 5 1312-1315 *KP HC2 FIFO error. cleared. MCB 100 3 1342-1345 *PT HC2 FIFO error. cleared. 21522 100 3 1348 Operator is A. Sowinski. 1352-1356 *LA rotation error. FRM 100 4 1352-1356 *FD rotation error. FRM 100 4 1411 BR reports wea: scattered low clouds, dry. 1436 PT reports wea: clear, calm and dry. 1503 OV reports wea: clear, dry. 1507-1510 *LA FRM rotation error. FRM 100 3 1524 KP reports wea: clear, dry and windy. 1825 %LA,PT Windy. 1922 MK reports wea: scattered cirrus clouds. 2110 %OV Windy. 2133 New Operator: Paul Padilla 2133-2322 %PT,KP occasional formatter errors, always during 21522 non-recording scans, so no downtime is being noted 2148 KP WEA: Clear and dry. 2215-2218 *KP AZ #1 armature over current, autostowed. ACU 100 3 Due to wind? 2249 OV WEA: cloudy 0000 Date is 2008OCT11/285. 0428 %KP windy, Az motor currents running a little high 0545 New Operator is Alan Kerr 0723-0757 *KP out of point. 55+ MPH wind stow. going WEA 100 34 to wait for the wind to lessen. 0742-0744 *PT HC2 FIFO. cleared. 21522 100 2 0758-0800 *PT HC2 FIFO. cleared. 21522 100 2 0843-0846 *KP out of point. 55+ wind stow. repointed. WEA 100 3 0935-1042 *KP 55+ wind stow. waiting again. WEA 100 67 1011-1043 *PT out of point. COM 100 32 1018 %PT precipitation starting. 1044-1047 *KP wind stow. WEA 100 3 1044 PT precipitation stopped. 1112-1114 *KP wind stow. WEA 100 2 1123-1206 *KP wind stow. 60+ MPH (25+ M/S). leaving it WEA 100 43 in stow for now. 1352 New operator is Betty Ragan. 1409 LA reports broken low clouds, dry conditions. 1420 %KP reports overcast, fog, light mist, wet. 1500 %PT raining. 1535 OV reports cloudy, dry weather. 1600 %KP raining. 1604-1606 *PT autostowed. Power glitch. COM 100 2 1730 %LA webcam showing dark clouds. Tsys are high. 1851 MK reports clear weather. 1900 FD now showing an incorrect temperature reading 23627 of -9.9C. Actual temperature is more like 25C. Dew point reading is correct, according to site tech. 1907 %LA raining. 1919-1922 *OV autostowed due to wind. WEA 100 3 1934-1936 *OV wind stow again. WEA 100 2 1951-1954 *OV stowed. WEA 100 3 2014-2030 *OV stowed again. Will leave it out for a little WEA 100 16 while. 2032-2034 *PT stowed. Power glitch. COM 100 2 2052-2053 *PT out of point due to power glitch again. COM 100 1 2144-2145 *OV stowed again. WEA 100 1 2207-2210 *OV wind stow. WEA 100 3 2224-2226 *PT stowed. Power glitch. COM 100 2 2257-2258 *PT stowed. Power glitch. COM 100 1 2322 End. Downtime Summary: Total downtime : 283 min Percentage downtime of observing: 1.1% Average downtime per hour : 0.7 min Total scheduled observing time (# Antennas): 25168 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard. VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C082B Observer: Krichbaum, T. Project type: VLBI Obs filename: c082bcrd.* Date/Day: 2008OCT11/285 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Betty Ragan Vclock, gps delta, UT1-UTC all check OK. 2327 Begin at all sites except MK. 2327-1100 %FD air temperature reading is incorrect. 23627 2327 Windy at LA, PT, KP, OV and MK. 2327 *PT raining. 0000 Date is 2008OCT12/286. 0121 New operator: T. Perreault 0211 MK joins 1100 FD temperature reading is correct again. 1338 New operator is Betty Ragan. 1716 *FD raining. 0000 Date is 2008OCT13/287. 0113-0115 *PT stowed. Probably a power glitch. COM 100 2 0115-0117 *OV stowed. Power glitch. COM 100 2 0132 New operator: T. Perreault 0434-0439 *OV autostowed. Power glitch. Repointed. COM 100 5 1245 FD wea: clear, no precipitation 1336 New operator is Betty Ragan. 1407 LA reports clear, dry weather. 1411 BR reports scattered low clouds, dry dish. 1434 OV weather: clear and dry. 1449 KP weather: clear and dry. 1602-1603 *PT stowed. Power glitch. COM 100 1 1845-1847 *OV stowed. Power glitch. COM 100 2 1851 MK reports clear weather. 2145 New operator is Heidi Gerhardt 2156-2158 %PT comm down, still observing COMM 0000 Date is 2008OCT14/288. 0021 *NL raining 0031-0309 %PT comm down, still observing (intermittent) COMM on dial back up now 0443-0448 %Comm down at all the sites, router is being reset due to internal problems. Should be back up shortly and should not affect data or observing. 0546 New Operator: Paul Padilla 0546 *NL Precipitation 0728 *FD Precipitation 1256 *FD WEA: Overcast, damp, light rain. 1317 *LA Precipitation 1351 New operator is Betty Ragan. 1417 BR reports a few high scattered clouds; dish is dry. 1431 OV weather: clear and dry. 1510-1513 *KP stowed. Power glitch. COM 100 3 1517 KP weather: clear and dry. 1525 FD reports overcast, intermittent light rain. 1921 MK weather: clear. 2013-2015 *PT stowed. Power glitch. COM 100 2 2147 New operator is Heidi Gerhardt 0000 Date is 2008OCT15/289. 0207 %OV 35+ mph wind caution 0252 %KP 35+ mph wind caution 0509-0510 %BR comm down, still observing COMM 0538 New Operator: Paul Padilla 0540-0543 *PT autostowed, power glitch. COM 100 3 0611 %FD Precipitation 0845 %NL Precipitation 1132 %FD WEA: Fog 1144 *NL Precipitation 1253 *FD WEA: Overcast, Light Rain 1327 PT WEA: Clear, Calm, Dry, and Cold 1339 New operator is Jim Ogle. 1416 BR reports a few high scattered clouds, frost on dish. 1505 OV reports clear skies. 1525 BR reports clear skies. 1653 End. Downtime Summary: Total downtime : 20 min Percentage downtime of observing: 0.0% Average downtime per hour : 0.0 min Total scheduled observing time (# Antennas): 42928 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.